A train normally travels 60 miles at a certain speed. one day,due to bad weather, the train's speed is reduced by 10mph so that the journey takes 3 hours longer. find the normal speed?
Answer by checkley77(12844)

D=RT
60=RT
R=60/T
60=(R-10)(T+3)
60=(60/T-10)(T+3)
60=60T/T+180/T-10T-30
60=60+180/T-10T-30
60-60+30=180/T-10T
30=(180-10T^2)/T
30T=180-10T^2
10T^2+30T-180=0
10(T^2+3T-18)=0
10(T+6)(T-3)
T-3=0
T=3
R=60/3
R=20 MPH
PROOF:
60=(20-10)(3+3)
60=10*6
60=60